OpenStudy (anonymous):

Y= 1/2X+2 X-2Y=-4 WRITE THE SOLUTION AS AN ORDERED PAIR

Parth (parthkohli):

So,\[\rm x - 2\left({1 \over 2}x + 2\right) = -4\]

OpenStudy (anonymous):

X-X-4=-4....

Parth (parthkohli):

Yes.

Parth (parthkohli):

lol wait a sec

Parth (parthkohli):

\[\rm 2y = x +4\]

Parth (parthkohli):

Hmm...

Parth (parthkohli):

Infinitely many solutions.

OpenStudy (anonymous):

HOW DID U GET THE 2Y

OpenStudy (asnaseer):

when you have two equations in two unknowns, then you can think of each equation as representing a line and the solution represents where these two lines intersect, e.g.:  However, if the two lines are parallel, then they will never intersect, and hence there will be no solution. Similarly, if the two lines represent the SAME line, then they will overlap each other and there will be an infinite number of solutions.